OpenStudy (anonymous):

Simplifying Radicals I need to know how to simplify radicals. Could someone go through the process? Use 5 radical 128 + 2 radical 18

OpenStudy (anonymous):

Is this the equation? \[5\sqrt{128}+2\sqrt{18}\]

OpenStudy (anonymous):

Yes, just when posting a question I don't get all the tools.

OpenStudy (anonymous):

Let's do 2 radical 18 first. What we are looking for a square values of 18 or factors of 18/ So what times what equals 18?

OpenStudy (anonymous):

2x9 which \[\sqrt{9} = 3\] so \[3\sqrt{2}\]

OpenStudy (anonymous):

Great but you need to multiply the 3 by 2 so you get 6 rational 2. Do you see this?

OpenStudy (anonymous):

Oh yeah.

OpenStudy (anonymous):

Ok now what about the other radical?

OpenStudy (anonymous):

So right now I have \[5\sqrt{128}+6\sqrt{2}\]

OpenStudy (anonymous):

\[5\sqrt{64*2}\]

OpenStudy (anonymous):

\[40\sqrt{2}\]

OpenStudy (anonymous):

So, \[40\sqrt{2}+6\sqrt{2}=46\sqrt{2}\]

OpenStudy (anonymous):

I think I was confused because I had \[6\sqrt{2}\] wrong.

OpenStudy (anonymous):

Great job....

OpenStudy (anonymous):

So, basically I need to know mt times tables real well?

OpenStudy (anonymous):

Yes when adding or subtracting you are looking to find "like radicals".